OpenStudy (anonymous):

Simplify the expression: 14/(3x) + (x+5)/(3x) so i tried doing this and i got x+19 over 3x

OpenStudy (astrophysics):

\[\frac{ 14 }{ 3x } + \frac{ (x+5) }{ 3x }\] this correct

OpenStudy (astrophysics):

They both have a common denominator of 3x so you can just add it regularly

OpenStudy (astrophysics):

\[\frac{ 14 }{ 3x } + \frac{ (x+5) }{ 3x } = \frac{ 14+(x+5) }{ 3x } = \frac{ 19+x }{ 3x }\]

OpenStudy (anonymous):

14 + x+ 5

OpenStudy (astrophysics):

You can leave it like this

OpenStudy (anonymous):

Oh the little / is actually for divide thats how my school rights it

OpenStudy (anonymous):

we got that you are good

OpenStudy (astrophysics):

Or you can write it as \[\frac{ 19 }{ 3x }+\frac{ 1 }{ 3 }\] it's the same thing